# Stochastic Calculus for Pathwise Observables of Markov-Jump Processes: Unification of Diffusion and Jump Dynamics

### v1

@@ A.10 Optimization of CTUR
Figure 14: Shown are quality factor of the CTUR in (a-c) using values of $c(t)$ shown in (d-f) in the calmodulin model. The values of $c(t)=c^{*}(t)+\frac{\gamma}{5}(c^{*}(t)-c^{\dagger}(t))$ are shown for values $10\leq\gamma\leq 10$, where $\gamma=-10$ is black and $\gamma=10$ is yellow. The initial condition is $p_{i}(0)=\delta_{i6}$ and the density uses $V_{i}(\tau)=\delta_{i3}$ in each panel. The transition weights are $\bm{\kappa}^{(12)}$ in (a) and (d), $\bm{\kappa}^{(56)}$ in (b) and (e), and $\bm{\kappa}^{(26)}$ in (c) and (f). The most optimal $c(t)=c^{*}(t)$ is marked by diamonds and the least optimal $c(t)=c^{\dagger}(t)$ is marked by squares. In (a-c) the $\mathcal{Q}$ for $c=0$ is included as a blue line.
In Fig. 14 we further elaborate on the influence of $c(t)$ in the CTUR. The quality factor $\mathcal{Q}$ for the calmodulin system (see Fig. 6) for currents with $\bm{\kappa}^{(12)}$, $\bm{\kappa}^{(56)}$, and $\bm{\kappa}^{(26)}$ (recall the notation $\kappa_{ij}^{(kl)}=\delta_{ik}\delta_{jl}-\delta_{il}\delta_{jk}$) and density with $V_{i}(\tau)=\delta_{i3}$ is shown in Fig. 14a-Fig. 14c. Various values of $c(t)$ are used, including $c(t)=0$, $c^{\dagger}(t)$, and $c^{*}(t)$. The absolute values of the considered $c(t)$ are shown in Fig. 14d-Fig. 14f. As can be easily seen, the behavior of $c^{\dagger}(t)$ and $c^{*}(t)$ is highly nontrivial and changes greatly in magnitude.

### v2

@@ A.10 Optimization of CTUR
Figure 14: Shown are quality factor of the CTUR in (a-c) using values of $c(t)$ shown in (d-f) in the calmodulin model. The values of $c(t)=c^{{\color[rgb]{0,0,0}\text{opt}}}(t)+\frac{\gamma}{5}(c^{{\color[rgb]{0,0,0}\text{opt}}}(t)-c^{{\color[rgb]{0,0,0}\min}}(t))$ are shown for values $10\leq\gamma\leq 10$, where $\gamma=-10$ is black and $\gamma=10$ is yellow. The initial condition is $p_{i}(0)=\delta_{i6}$ and the density uses $V_{i}(\tau)=\delta_{i3}$ in each panel. The transition weights are $\bm{\kappa}^{(12)}$ in (a) and (d), $\bm{\kappa}^{(56)}$ in (b) and (e), and $\bm{\kappa}^{(26)}$ in (c) and (f). The most optimal $c(t)=c^{{\color[rgb]{0,0,0}\mathrm{opt}}}(t)$ is marked by diamonds and the least optimal $c(t)=c^{{\color[rgb]{0,0,0}\min}}(t)$ is marked by squares. In (a-c) the $\mathcal{Q}$ for $c=0$ is included as a blue line.
In Fig. 14 we further elaborate on the influence of $c(t)$ in the CTUR. The quality factor $\mathcal{Q}$ for the calmodulin system (see Fig. 6) for currents with $\bm{\kappa}^{(12)}$, $\bm{\kappa}^{(56)}$, and $\bm{\kappa}^{(26)}$ (recall the notation $\kappa_{ij}^{(kl)}=\delta_{ik}\delta_{jl}-\delta_{il}\delta_{jk}$) and density with $V_{i}(\tau)=\delta_{i3}$ is shown in Fig. 14a-Fig. 14c. Various values of $c(t)$ are used, including $c(t)=0$, $c^{{\color[rgb]{0,0,0}\mathrm{opt}}}(t)$, and $c^{{\color[rgb]{0,0,0}\mathrm{min}}}(t)$. The absolute values of the considered $c(t)$ are shown in Fig. 14d-Fig. 14f. As can be easily seen, the behavior of $c^{{\color[rgb]{0,0,0}\mathrm{opt}}}(t)$ and $c^{{\color[rgb]{0,0,0}\mathrm{min}}}(t)$ is highly nontrivial and changes greatly in magnitude.